OpenStudy (anonymous):

can someone help in this? If 4x^2=25(y^2+2), then 50/(2x-5y) =?

OpenStudy (anonymous):

4x^2=25(y^2+2) 4x^2-25y^2-50 = 0 But 4x^2-25y^2 = (2x-5y)(2x+5y) (2x-5y)(2x+5y) = 50 => 50/(2x-5y) = (2x+5y)

OpenStudy (anonymous):

good explanation giorgia ver gud

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

let me think again

OpenStudy (anonymous):

http://twiddla.com/579589

OpenStudy (anonymous):

can u answer me there

OpenStudy (anonymous):

r u here